OpenStudy (anonymous):

10c-f=-13+cd solve for c. show work please

OpenStudy (debbieg):

Welcome to Open Study! You'll first want to get both of the terms that involve c on the same side, and all the non-c terms on the other side. Any ideas on how to do that?

OpenStudy (anonymous):

Ok first you need to get the variables to one side. you should get 10c-f=-13+cd =10c-cd-f=-13 =10c-cd=f-13 =c(10-d)=f-13 =c=(-13+f)/(10-d) <--- your answer

OpenStudy (anonymous):

thanks man